OpenStudy (anonymous):

a 300-n force acts a 25-kg object calculate the acceleration of the object

OpenStudy (loser66):

F=ma ---> a = F/m = 300N/ 25 kg = 12m/s^2

OpenStudy (anonymous):

thank you for the answer